Control limits for np chart if a standard value of P is available:
1
Expert's answer
2021-09-21T11:04:21-0400

Solution:

Then control limits for np chart if a standard value of P is available can be calculated as:

"UCL=np+3\sqrt{np(1-p)}
\\LCL=np-3\sqrt{np(1-p)}"